### v1

# Watermark Anything with Localized Messages

Abstract
Image watermarking methods are not tailored to handle small watermarked areas. This restricts applications in real-world scenarios where parts of the image may come from different sources or have been edited. We introduce a deep-learning model for localized image watermarking, dubbed the Watermark Anything Model (WAM). The WAM embedder imperceptibly modifies the input image, while the extractor segments the received image into watermarked and non-watermarked areas and recovers one or several hidden messages from the areas found to be watermarked. The models are jointly trained at low resolution and without perceptual constraints, then post-trained for imperceptibility and multiple watermarks. Experiments show that WAM is competitive with state-of-the art methods in terms of imperceptibility and robustness, especially against inpainting and splicing, even on high-resolution images. Moreover, it offers new capabilities: WAM can locate watermarked areas in spliced images and extract distinct 32-bit messages with less than 1 bit error from multiple small regions – no larger than 10% of the image surface – even for small $256\times 256$ images.
[…]
[Code and models]https://github.com/facebookresearch/watermark-anything
1 Introduction
Invisible image watermarking embeds information into image pixels in a way that is imperceptible to the human eye and yet robust. It was initially developed for intellectual property and copy protection, such as by Hollywood studios for DVDs. However, the applications of watermarking are evolving, particularly in light of the recent development of generative AI models (Kušen and Strembeck, 2018). Regulatory acts such as the White House executive order (USA, 2023), the Californian bill, the EU AI Act (Parliament and Council, 2024), and Chinese AI governance rules (of the People’s Republic of China, 2023) require AI-generated content to be easily identifiable. They all cite watermarking as either compulsory or a recommended measure to detect and label AI-generated images.
Image splicing is one of the most common manipulations, whether applied for benign or malicious purposes (Christlein et al., 2012; Tralic et al., 2013). Splicing involves adding text or memes on a large portion of the image or extracting parts of images and overlaying them on others (Douze et al., 2021). It can bypass the state-of-the-art watermarking techniques, which take one global decision per image under scrutiny. Indeed, in traditional watermarking, the watermark signal fades away and is no longer detected as the surface of the watermarked area decreases. Besides, these techniques poorly answer the paradoxical question of deciding whether an image should be considered watermarked if only a small part carries the watermark. A positive decision triggered by a small area might be unfair to artists who use AI models for inpainting or outpainting. On the other hand, not being robust enough to splicing opens the door to easy removal.
Figure 1: Overview. (a) The embedder creates an imperceptible image modification. (b) Traditional transformations (cropping, JPEG compression, etc.) and/or advanced manipulations (mixing watermarked and non-watermarked images, inpainting, etc.) may be applied to the image. (c) The extraction creates a segmentation map of watermarked parts and retrieves one or several messages.
To address these issues, this paper redefines watermarking as a segmentation task, giving birth to the Watermark Anything Models (WAM). Our motivation is to disentangle the strength of the watermark signal from its pixel surface, in contrast to traditional watermarking. More precisely, the WAM extractor detects if the watermark is present and extracts a binary string for every pixel rather than predicting a message for the whole image. These outputs are post-processed according to the final task. For global detection, the image is deemed watermarked if the proportion of watermarked pixels exceeds a user-defined threshold. For global decoding, a majority vote recovers the hidden message. A new application, out of the reach of traditional robust watermarking, is the localization of watermarked areas and the extraction of multiple hidden messages. For that purpose, we choose to apply the DBSCAN clustering algorithm over the pixel-level binary strings because it does not require any prior on the number of watermarks (or centroids). This is detailed in Sec. 3.
These new functionalities require a training with new objectives, that is split into two phases. The first phase pre-trains the embedder and extractor models for low-resolution images. It essentially targets the robustness criterion. The embedder encodes a $n_{\text{bits}}$-bit message into a watermark signal that is added to the original image. The augmenter randomly masks the watermark in parts of the image and augments the result with common processing techniques (e.g., cropping, resizing, compression). The extractor then outputs a ($1+n_{\text{bits}}$)-dimensional vector per pixel to predict the parts of the image that are watermarked and decode the corresponding messages. Detection and decoding losses are used as training objectives. The second training phase targets the following new objectives: (1) minimize the watermark’s visibility in alignment with the human visual system, (2) allow for multiple messages within the same image. This two-stage training is less prone to instability, compared to previous use of adversarial networks and divergent objectives (Zhu et al., 2018). It also trains the extractor on both watermarked and non-watermarked images, for the first time in the literature. This increases the performance and the robustness of the detection.
We first compare WAM with state-of-the-art methods for regular tasks of watermark detection and decoding on low and high-resolution images. Our results show that WAM achieves competitive performance in terms of imperceptibility and robustness. To further highlight the advantages of WAM, we then evaluate its performance on tasks that are not considered in the literature. Namely, we evaluate the localization accuracy between the predicted watermarked areas and the original mask and assess the ability to detect and decode multiple watermarks in a single image. For instance, when hiding five 32-bit messages, each in a 10% area of the image, detection of watermarked areas achieves more than 85% mIoU, even after images are horizontally flipped and the contrast adjusted, and bit accuracy (for a total of 160 bits) achieves more than 95% under the same augmentation (Sec. 5.5).
In summary, our contributions are:
  • the definition of watermarking as a segmentation task;
  • a two-stage training able to strike a good trade-off between invisibility and robustness even for multiple watermarks and high-resolution images;
  • WAM, an embedder/extractor model competitive with state-of-the-art methods;
  • the highlight of new capabilities, localization of watermarks and extraction of multiple messages as depicted in Fig. 1, together with specially designed evaluations.
2 Related Work
Semantic segmentation
aims to predict a category label for every pixel in an image. FCN (Long et al., 2015) employs a convolutional network predicting pixel-wise classification logits. More recent works (Zheng et al., 2021; Strudel et al., 2021) aggregate hierarchical context and are based on a ViT encoder (Dosovitskiy, 2020), followed by a decoder that generates the pixel-level predictions – similar to our extractor’s architecture. On the other hand, instance segmentation identifies individual objects within an image (He et al., 2017; Carion et al., 2020). Recent literature, such as MaskFormer (Cheng et al., 2021) and Mask2former (Cheng et al., 2022), unifies semantic and instance segmentation. Segment Anything (Kirillov et al., 2023) and follow-ups (Zhang et al., 2023; Zhao et al., 2023a; Ke et al., 2024; Ma et al., 2024; Ravi et al., 2024) go one step further. They now allow users to segment any object in images or videos using prompts, which are points, bounding boxes, or natural language.
Robust watermarking
was first designed for copyright protection (Cox et al., 2007). Traditional methods operate either in the spatial domain (Van Schyndel et al., 1994; Nikolaidis and Pitas, 1998) or in the frequency domain modifying components through an invertible transform (e.g., Fourier, Wavelet) (Cox et al., 1997; Urvoy et al., 2014). These methods were progressively replaced by deep-learning ones which remove the expert need to crafting the transforms such that the embedding is imperceptible and robust. The first approach (Vukotić et al., 2018, 2020; Fernandez et al., 2022; Kishore et al., 2022; Chen et al., 2022) embeds the watermark into the representations of an off-the-shelf pre-trained model. Our method is inspired by the second approach, which jointly trains deep learning-based architectures to embed and extract watermarks while being robust to augmentations seen during training, such as HiDDeN (Zhu et al., 2018) and followers (Zhang et al., 2019b; Luo et al., 2020; Tancik et al., 2020; Ma et al., 2022; Bui et al., 2023b, a; Pan et al., 2024).
The two tasks of watermarking, detection and decoding, are rarely performed together. Detection of the watermark, a.k.a., zero-bit watermarking, distinguishes watermarked content from original images. Decoding of a hidden message, a.k.a., multi-bit watermarking, implicitly assumes that the content under scrutiny is watermarked. One possibility for combining the two tasks is to reserve $n_{\text{det}}$ bits of the message as a detection segment – that should match a fixed pattern – with the remaining bits carrying the actual payload. Assuming that the message decoded from a non-watermarked image is random, the False Positive Rate equals $2^{-n_{\text{det}}}$. However, a recent study (Fernandez et al., 2023a, App. B.5) shows that the decoded bits are neither equiprobable nor independent. In other words, it is difficult to decode a hidden message while being confident in detecting a watermark.
Very few papers in the literature deal with watermarking objects in picture (Bas et al., 2001; Barni et al., 2005). This trend was abandoned together with the object-oriented MPEG-4 video codec.
Active tamper localization
relies on semi-fragile watermarking. Areas where the watermark is not recovered are deemed tampered. This idea appears early in the literature (Kundur and Hatzinakos, 1999; Lin et al., 2000), although deep-learning methods offer significant improvements (Asnani et al., 2022). There is a trade-off between the granularity of the localization of the tampered areas and the semi-fragility of the watermark. The biggest difficulty is to design a watermark robust to benign transformations (e.g., image compression) but fragile to malicious editing.
Our method combines detection and decoding together with a precise segmentation of the watermarked areas. This natively answers applications ranging from copyright protection, detection of AI-generated objects in images and tampering localization. EditGuard (Zhang et al., 2024) sequentially embeds a fragile watermark for localization and a robust watermark for copyright protection. Although the approach shares similarities with ours, it is not robust to geometric augmentations such as cropping or perspective changes. In contrast, our approach is conceptually simpler since a single embedding is used for detection and decoding. It offers better robustness, and enables the extraction of multiple watermarks. The approach most similar to ours is AudioSeal (San Roman et al., 2024), which introduces localized watermarking in the audio domain, but does not handle multiple messages.
An extended related work with more details on the methods mentioned above is presented in App. 8.
3 Detection, Localization, and Message Extraction
Before introducing our method and its training, this section presents several applications, i.e., how to use WAM’s extractor outputs for watermark localization, zero-bit detection, and decoding of multiple messages within the same image.
The key feature is the extractor $\mathrm{ext}_{\theta^{*}}$ ($\theta^{*}$ refers to the weights of the model after training) which outputs a tensor $y=\mathrm{ext}_{\theta^{*}}(x)$ of dimensions $(1+n_{\text{bits}},h,w)$ for an input image $x\in\mathbb{R}^{3\times h\times w}$. This tensor consists of a watermark detection mask $y^{\text{det}}\in[0,1]^{1\times h\times w}$ and a decoding mask $y^{\text{dec}}\in[0,1]^{n_{\text{bits}}\times h\times w}$. We denote by $y^{\text{det}}_{i}\in[0,1]$ the predicted detection score for pixel $i$, and by $y^{\text{dec}}_{i}=[y^{\text{dec}}_{i,1}\dots y^{\text{dec}}_{i,n_{\text{bits}}}]\in[0,1]^{n_{\text{bits}}}$ its decoded message.
Localization
(or pixel-level detection) spots watermarked pixels of the image. A pixel is deemed watermarked if its detection score $y^{\text{det}}_{i}$ exceeds a threshold $\tau$. Typically, we set $\tau$ empirically, by measuring the False Positive Rate (FPR) on all the pixels of a held-out training set (the FPR is the probability of a non-watermarked pixel being flagged as such).
Detection
(or image-level detection) decides if an image is globally watermarked. From the pixel-level detection score, an image soft detection score can naturally be computed as:
$s_{\text{det}}:=\frac{1}{h\times w}\sum_{i=1}^{h\times w}\mathbbm{1}\{y^{\text{det}}_{i}>\tau\}\in[0,1].$ (1)
The image is flagged if $s_{\text{det}}$ is higher than a threshold that is the proportion of watermarked pixels that the user considers enough to deem a content as watermarked.
Decoding
of a single message within an image is done by the following weighted average of the pixel-level soft predictions of the hidden message:
$\hat{m}_{k}=\left\{\begin{array}[]{ll}1&\text{if }\left[\frac{1}{\sum_{i}\mathbbm{1}\{y^{\text{det}}_{i}>\tau\}}\left(\sum_{i=1}^{h\times w}\mathbbm{1}\{y^{\text{det}}_{i}>\tau\}\cdot y^{\text{dec}}_{i,k}\right)\right]>0.5\,,\\[8.0pt] 0&\text{otherwise}\,.\end{array}\right.$ (2)
Decoding of multiple watermarks within one image uses a hard detection approach instead of the previous soft weighting. We isolate pixels detected as watermarked, i.e., pixels $i$ with $y^{\text{det}}_{i}>\tau$, and we compute the local decoded message $\tilde{m}_{i}$ such that for any $k\in\{1,\ldots,n_{\text{bits}}\}$, $\tilde{m}_{i,k}=\mathbbm{1}\{y^{\text{dec}}_{i,k}>0.5\}$. The DBSCAN (Density-Based Spatial Clustering of Applications with Noise) algorithm (Ester et al., 1996; Schubert et al., 2017) clusters the set of locally decoded messages. It outputs some centroids and an assignment for every pixel detected as watermarked. DBSCAN selects these centroids among the initial set, thus ensuring that the final decoded messages are binary words.
DBSCAN offers the advantage of not requiring a pre-defined number of clusters. Instead, we need to specify two parameters: (1) $\text{min}_{\text{samples}}$, the minimum number of points for a cluster to be considered valid, and (2) $\varepsilon$, the maximum distance between two samples for one to be considered as in the neighborhood of the other. For further details on the DBSCAN algorithm, see App. 9.1.
Handling multiple messages in a single image makes WAM robust against attacks that involve splicing several watermarked images together, which can compromise the decoding of traditional watermarking schemes in real-world scenarios. It also enables active object detection (Asnani et al., 2024), where the objective is to track watermarked objects. Additionally, it allows for the identification of the use of multiple watermarked AI tools.
4 Watermark Anything Models
4.1 The model
WAM considers two joint models: a watermark embedder $\mathrm{emb}_{\theta}$ and a watermark extractor $\mathrm{ext}_{\theta}$, where $\theta$ gathers the parameters of these functions. The embedder defines the watermark procedure for hiding a message into an image, while the extractor detects watermarking and decodes messages. This section discusses our choices for the different pieces that constitute WAM. App. 10.1 gives the technical details of the network architectures.
Figure 2: The first training phase of WAM, as described in Sec. 4.2, jointly trains the watermark embedder to predict an additive watermark and the watermark extractor to detect watermarked pixels and decode the hidden message. In between, the augmenter 1) splices the watermarked and the original images based on a random mask and 2) applies classical image transformations.
The watermark embedder
($\mathrm{emb}_{\theta}$) consists of an encoder, a binary message lookup table, and a decoder. The encoder represents an image in a latent space. The lookup table is used to translate the message into a tensor which is concatenated to the image representation. From that, the decoder outputs a watermark signal, which is added to the original image with adequate scaling.
The autoencoder that we consider is based on the architecture of the variational autoencoder of LDM (Rombach et al., 2022). Its encoder, $\mathrm{enc}_{\theta}$, compresses a $h\times w$ input image $x$ into a latent $z\in\mathbb{R}^{d_{\text{z}}\times h^{\prime}\times w^{\prime}}$ (with downsampling factor $f=h/h^{\prime}=w/w^{\prime}$). The binary message lookup table $\mathcal{T}_{\theta}$ is of shape $(n_{\text{bits}},2,d_{\text{msg}})$. Each bit of the message $m$ is mapped to the embedding $\mathcal{T}_{\theta}(k,m_{k},\cdot)\in\mathbb{R}^{d_{\text{msg}}}$, depending on its position $k\in\{1,\ldots,n_{\text{bits}}\}$ and its value $m_{k}\in\{0,1\}$. The $n_{\text{bits}}$ embeddings are averaged, resulting in a vector of size $d_{\text{msg}}$ which is repeated to form a tensor of shape $(d_{\text{msg}},h^{\prime},w^{\prime})$. The concatenation to the image representation yields an activation of shape $(d_{\text{z}}+d_{\text{msg}})\times h^{\prime}\times w^{\prime}$. The decoder, $\mathrm{dec}_{\theta}$, maps back this activation to the watermark signal $\delta_{\theta}(x,m)$ of the same shape as the original image. The range of this signal is $[-1,1]$ because the last layer of $\mathrm{dec}_{\theta}$ is a hyperbolic tangent activation. It is finally added to the original image to produce the watermarked output $x_{\text{m}}=\mathrm{emb}_{\theta}(x)=x+\alpha\cdot\delta_{\theta}(x,m)$. Parameter $\alpha\in\mathbb{R}_{+}$ is called the watermark strength as it controls the distortion applied to the original image.
The watermark extractor
($\mathrm{ext}_{\theta}$) takes on the dual role of detecting watermarked pixels and decoding their embedded messages. We use an architecture similar to SETR (Zheng et al., 2021) and Segment Anything (Kirillov et al., 2023). It comprises a ViT encoder paired with a pixel decoder that upsamples the embeddings to the original image size. The watermark extractor outputs a vector of size $1+n_{\text{bits}}$ for every pixel, as described in Sec. 3.
High-resolution.
WAM operates at a fixed resolution of $h\times w$. To extend it to higher resolutions, an anisotropic scaling resizes the image to $h\times w$ and $\mathrm{emb}_{\theta}$ computes the watermark signal $\delta$ from this resized image. A bilinear interpolation scales $\delta$ back to the size of the original image. The extraction also operates at $h\times w$, by resizing all images before feeding them to the network. Therefore the extraction process remains consistent with the pre-training conditions. We only use these interpolations at embedding time. WAM is thus only trained on low resolution images (Sec. 4.2) but can be used for high-resolution images too, which represents an important training compute gain. A similar approach also appears in the recent literature (Bui et al., 2023a).
4.2 Pre-training models for localized message embedding and extraction
The objective of this first training phase is to obtain a robust and localizable watermark hiding a $n_{\text{bits}}$-bit message inside parts of an image, without caring much about imperceptibility. This stage does not incorporate any perceptual loss: our goal is to achieve perfect localization and decoding even after severe augmentations. Figure 2 illustrates the detailed process.
Augmentations
are two-step processes. They first splice images $x$ and $x_{\text{m}}$ based on a binary mask and then apply usual image transformations to improve robustness.
The first step randomly samples a mask among full masks, rectangles, irregular shapes, or object segmentation maps (possibly provided by the training dataset). These masks are inverted with probability $0.5$. It yields a mask $r_{\text{mask}}\in[0,1]^{h\times w}$. Figure 2 shows examples of masks, with more samples displayed in Figure 6 of App. 10.2. The first step then computes the spliced image as $x_{\text{masked}}=r_{\text{mask}}\odot x_{\text{m}}+(1-r_{\text{mask}})\odot x$, where $\odot$ denotes component-wise multiplication.
The second step of the augmentation applies a processing typical from real-world image editing including geometric transformations (identity, resize, crop, rotate, horizontal flip, perspective) and valuemetric adjustments (JPEG, Gaussian blur, median filter, brightness, contrast, saturation, hue). The geometric transformation alters the positions of watermarked pixels, so it is also applied to the mask to keep track of the watermarked areas. This adjusted mask is the ground truth for the localization and it is denoted $y^{\text{det},\star}$ in the sequel. These augmentations are relatively strong to reflect the transformations found on media sharing websites, as opposed to professional photo publications. Additional details and discussions on augmentations are postponed to App. 10.4.
When a given image goes through the embedder-augmentation-extractor sequence, the final output depends on the models’ parameters $\theta$ (omitted so far for clarity):
$y(\theta)=[y^{\text{det}}(\theta),y^{\text{dec}}(\theta)]=\mathrm{ext}_{\theta}\ \Big{(}\ \text{transformation}\ \big{(}\ \text{masking}\ \left(\ \mathrm{emb}_{\theta}(x,m),\,x\ \right)\ \ \big{)}\ \ \Big{)}\ .$ (3)
Objectives.
The training minimizes the objective function $\ell(\theta)$ which is a linear combination of the detection and decoding losses: $\ell(\theta)=\lambda_{\text{det}}\cdot\ell_{\text{det}}(\theta)+\lambda_{\text{dec}}\cdot\ell_{\text{dec}}(\theta)$.
The detection loss is the average of the pixel-wise cross-entropy between $y^{\text{det}}(\theta)\in[0,1]^{h\times w}$ and the ground truth $y^{\text{det},\star}\in\{0,1\}^{h\times w}$ (pixel watermarked or not). Similarly, the decoding loss is the average of the pixel-wise and bit-wise binary cross-entropy between $y^{\text{dec}}_{i,k}(\theta)$ and $m_{k}$ only over the watermarked pixels, where $m$ is a random message originally embedded in that image. For a given image and message, $\ell_{\text{det}}$ and $\ell_{\text{dec}}$ are:
$\ell_{\text{det}}(\theta)$ $=\frac{-1}{h\times w}\sum_{i=1}^{h\times w}\left[y^{\text{det},\star}_{i}\log(y^{\text{det}}_{i}(\theta))+(1-y^{\text{det},\star}_{i})\log(1-y^{\text{det}}_{i}(\theta))\right],$ (4)
$\ell_{\text{dec}}(\theta)$ $=\frac{-1}{n_{\text{bits}}\times\sum_{i=1}^{h\times w}y^{\text{det},\star}_{i}}\sum_{i=1}^{h\times w}y^{\text{det},\star}_{i}\sum_{k=1}^{n_{\text{bits}}}\left[m_{k}\log(y^{\text{dec}}_{i,k}(\theta))+(1-m_{k})\log(1-y^{\text{dec}}_{i,k}(\theta))\right].$ (5)
4.3 Post-training for imperceptibility and multiple watermarks
The model trained in the first phase (Sec. 4.2) produces a too visible watermark. Moreover, it cannot deal within multiple watermarks in an image. The second training phase addresses these issues.
Perceptual heatmap.
(a) Without JND
(b) With JND
Figure 3: Impact of the JND map on imperceptibility. (Left) After the first training phase, the watermark is highly perceptible. (Right) When applying the JND attenuation, it is hidden in areas where the eye is not sensitive to changes, which makes it less visible. Fine-tuning with the JND recovers the initial robustness. The difference is displayed as $10\times\text{abs}(x_{\text{m}}-x)$.
The Just-Noticeable-Difference (JND) map is a hand-crafted model of the minimum artifact perceivable by a human at every pixel. For instance, artifacts on flat, uniform areas are more visible than on textured ones. The JND map was introduced by Chou and Li (1995) and later used by Wu et al. (2017). App. 9.2 gives more details on its computation.
Given an image $x$, we have $\text{JND}(x)\in\mathbb{R}^{3\times h\times w}$. We modulate the intensity of the watermark per pixel at embedding time: the final watermarked image is computed as $x_{\text{m}}=x+\alpha_{\mathrm{JND}}\cdot\mathrm{JND}(x)\odot\delta_{\theta}(x,m)$. In practice, we found that applying this JND on the model trained in the first phase slightly degrades the detection and decoding performance. The fine-tuning cancels this degradation. Applying JND only in the second training phase makes training easier than previous methods relying on “perceptual” or “contradictory” losses (like StegaStamp (Tancik et al., 2020) or HiDDeN (Zhu et al., 2018)).
Multiple watermarks.
Since the first training phase (Sec. 4.2) uses a single message, it tends to decode a constant message across pixels even on image regions with different messages. The second training phase introduces several masks to address this limitation. The masks are generated as before with distinct random messages in each of the masked areas. The number of disjoint masks goes from 1 to 3 with probabilities 0.6, 0.2, and 0.2 respectively (see App. 10.2).
The detection loss $\ell_{\text{det}}$ takes the union of all masks as ground truth $y^{\text{det},\star}_{i}$. The decoding loss $\ell_{\text{dec}}$ is computed separately for each message and the losses are summed up. The scenario of multiple watermarks within a single image arises when watermarked images are combined. Unlike other methods producing a global message, WAM can now distinguish and decode each message separately.
5 Experiments & Results
5.1 Implementation details
We provide here further implementation details using the notations introduced in Sec. 4.1, and the rest of the architectures are described in App. 10.1. All experiments are run with $n_{\text{bits}}=32$. There is a total of $1.1$M parameters for the embedder and $96$M for the extractor (equivalent to a ViT-base). In our experiments, we found it necessary to increase the size of the extractor compared to the embedder. This is probably because the extractor has more tasks to handle, such as segmenting and extracting multiple messages, while the embedder only needs to embed one message into an image. Also, it is important for the embedding process to be quick since it happens on the user-side, so we keep it voluntarily small to be fast and usable at scale. We train our model on the MS-COCO training set with blurred faces (Lin et al., 2014), that contains 118,000 images, many of them with segmentation masks. We train at resolution $h\times w=256\times 256$, and with $f=8$ (so $h^{\prime}\times w^{\prime}=32\times 32$). The first training phase (Sec. 4.2) is optimized with AdamW (Kingma, 2014; Loshchilov, 2017) with a linear warmup of the learning rate in $5$ epochs from $1\times 10^{-6}$ to $1\times 10^{-4}$ and a cosine annealing to $1\times 10^{-6}$. We set $\lambda_{\text{dec}}=10$, $\lambda_{\text{det}}=1$, $\alpha=0.3$. We train with a batch size of $16$ per GPU for 300 epochs using 8 V100 GPUs which takes roughly 2 days. The second training phase (Sec. 4.3) further trains the model with the JND attenuation for 200 epochs, hiding up to 3 messages per image using either randomly sampled rectangles or segmentation masks. During this phase, $\alpha_{\text{JND}}=2$. App. 10.2 provides details on the generation of masks used during both training phases. For the extraction of multiple messages, we use the Scikit-learn DBSCAN implementation (Pedregosa et al., 2011).
5.2 Quality
Table 1 evaluates quantitatively the difference between the watermarked and original images with PSNR, SSIM and LPIPS (Zhang et al., 2018) metrics. We adapt the methods so that they are comparable in terms of imperceptibility. We show qualitative examples for COCO and DIV2k images in App. 12. The perceptual quality is good: the JND map successfully modulates the watermark signal in areas where the eye is not sensitive. Nevertheless, repetitive and regular patterns can be observed in some images when the bright or textured areas are big enough (furs of animals for instance).
Table 1: Evaluation of the watermark imperceptibility. We report the PSNR, SSIM, and LPIPS between watermarked and original images of COCO (low/mid-resolution) and DIV2k (high-resolution).
HiDDeN DCTDWT SSL FNNS TrustMark WAM (ours)
COCO PSNR ($\uparrow$) 38.2 37.0 37.8 37.7 40.3 38.3
SSIM ($\uparrow$) 0.98 0.98 0.98 0.98 0.99 0.99
LPIPS ($\downarrow$) 0.05 0.02 0.07 0.06 0.01 0.04
DIV2K PSNR ($\uparrow$) 38.4 38.7 38.2 39.0 39.1 38.8
SSIM ($\uparrow$) 0.98 0.99 0.98 0.99 0.99 0.99
LPIPS ($\downarrow$) 0.07 0.03 0.11 0.04 0.01 0.03
5.3 Detection and decoding
We benchmark WAM’s performance against several watermarking methods: DCTDWT (Al-Haj, 2007), HiDDeN (Zhu et al., 2018), TrustMark (Bui et al., 2023a), SSL (Fernandez et al., 2022), FNNS (Kishore et al., 2022). We also compare WAM to generation-time watermarking methods (Fernandez et al., 2023a; Wen et al., 2023) in App. 11.1. Although this list is not exhaustive, it covers the main types of state-of-the-art methods. For HiDDeN, we use a model which hides 48 bits. FNNS, SSL, and DCTDWT can hide messages or arbitrary length, we choose to hide 48 bits as well. We use the first 16 bits for detection and 32 bits for message decoding. For detection, an image is flagged as watermarked if at most 1 bit is wrongly decoded. This corresponds to a theoretical False Positive Rate (FPR) of $2.6\times 10^{-4}$, i.e., $2.6$ out of 10,000 images are falsely flagged as watermarked on average. TrustMark (Bui et al., 2023a) directly outputs a boolean (watermarked or not). We use the version that hides $40$ bits, but only use the first 32 bits. For all evaluations, we apply the watermark embedding and extraction at the original image resolution as described in the high-resolution paragraph of Sec. 4.1. Finally, for WAM, an image is flagged if $s_{\text{det}}$ (Eq. 1) is higher than 0.07. This value empirically delivers an approximately similar FPR on the COCO validation set.
We evaluate the robustness against various transformations: geometric (flip, crop, perspective, and rotation), valuemetric (adjustments in brightness, hue, contrast or saturation, median or Gaussian filtering, and JPEG compression), splicing: scenarios where only 10% of the image is watermarked superimposed onto the original or an other background. We also evaluate the robustness against inpainting, using the LaMa (Suvorov et al., 2022) model to inpaint areas of the watermarked image specified by random or segmentation masks. For COCO, we use the union of all segmentation masks for each image which in average corresponds to 30% of the image, and for DIV2k we use a randomly generated mask (because there are no segmentation masks) that covers around 35% of the image. We give details on the compared baselines in App. 10.3 and on the transformations in App. 10.4.
2(a) presents the detection and decoding results for low/mid-resolution images, averaged over the first 10k images of the COCO validation set – the detailed results for every augmentation used to form the different groups are in App. 11.3. 2(b) does the same for high-resolution images from the DIV2k (Timofte et al., 2018) validation set. For both distributions, WAM is competitive and even shows improved robustness on classical geometric/valuemetric transformations. Most importantly, WAM performs better on splicing or inpainting where other methods fail to achieve more than 90% TPR and Bit acc. This is true even if WAM was not explicitly trained on high resolution images or to be robust against inpainting. We also show examples of WAM’s detection masks after inpainting in Fig. 10 of App. 11, where some modified parts of images are not detected as watermarked anymore.
Table 2: Detection and decoding after image editing (detailed in Sec. 10.4). We show the bit accuracy (Bit acc.) between the encoded and decoded messages, the proportion of images correctly deemed watermarked (TPR), and the proportion of non-watermarked images falsely detected as watermarked (FPR), in %. Since HiDDeN, DCTDWT, SSL and FNNS do not naturally provide a detection result, we hide 48 bits and reserve 16 bits for detection, and flag an image as watermarked if it has strictly less than two bits incorrectly decoded (these baselines are detailed in Sec. 5.3).
(a) On the first 10k validation images of COCO (low to mid resolution)
Augmentations
None Geometric Valuemetric Inpainting Splicing
Method FPR TPR Bit acc. TPR Bit acc. TPR Bit acc. TPR Bit acc. TPR Bit acc.
HiDDeN 0.08 76.9 95.5 31.2 80.1 48.4 87.2 44.7 88.7 0.9 68.0
DWTDCT 0.02 77.1 91.4 0.0 50.5 13.6 58.1 47.8 81.7 0.8 59.9
SSL 0.00 99.9 100.0 14.3 76.5 70.5 92.1 67.7 91.1 0.4 58.9
FNNS 0.10 99.6 99.9 62.4 86.6 82.1 93.9 89.4 97.3 38.7 88.5
TrustMark 2.88 99.8 99.9 36.3 71.4 90.1 98.2 39.4 83.2 83.2 57.1
WAM (ours) 0.04 100.0 100.0 99.3 91.8 100.0 99.9 97.9 99.2 100.0 95.3
(b) On the 100 validation images of DIV2k (high resolution)
Augmentations
None Geometric Valuemetric Inpainting Splicing
Method FP TPR Bit acc. TPR Bit acc. TPR Bit acc. TPR Bit acc. TPR Bit acc.
HiDDeN 0 72.0 95.8 33.5 81.3 53.3 88.1 53.3 88.1 1.5 70.2
DWTDCT 0 75.0 88.9 0.0 50.6 18.4 58.9 73.0 86.7 31.5 71.6
SSL 0 100.0 100.0 32.5 83.6 75.8 92.8 95.0 96.1 0.5 59.8
FNNS 0 97.0 99.9 63.5 87.2 79.8 93.9 94.0 99.2 48.5 89.5
TrustMark 5 100.0 100.0 33.1 70.5 87.4 97.9 0.0 72.1 1.5 57.3
WAM (ours) 0 100.0 99.9 96.1 89.0 100.0 99.9 100.0 99.8 99.5 94.2
5.4 Localization
WAM and EditGuard (Zhang et al., 2024) are the only methods that provide watermark localization. We therefore consider images at fixed resolution $512\times 512$ (unlike in Sec. 5.3) to align with the setup of EditGuard. We nevertheless observe similar results at different resolutions for WAM. We focus on the COCO validation set and splice centered watermarked rectangles of various sizes within the images to ensure a well-controlled experiment. We then either apply no transformation or crop the upper left part of the image (25%), which is then resized to the original size. By doing so, the proportion of watermarked pixels is still the same as in the spliced image before cropping, which allows us to evaluate the robustness of the localization. Figure 7 of App. 10.5 illustrates this protocol.
Figure 4 reports the mean Intersection over Union (mIoU) to evaluate the localization, as commonly done in the segmentation literature. Figure 4 evaluates the bit accuracy – through localization – following Eq. 2. We observe that WAM accurately predicts both classes, even after cropping and resizing, except when the watermarked area covers 95% of the image, in which case the extractor tends to classify all pixels as watermarked. In terms of bit accuracy, WAM recovers in average 31 out of 32 bits even when only 10% of the $256\times 256$ image is watermarked, and around 25 bits when only 10% of a 25% crop is watermarked (which corresponds to 2.5% of the overall number of pixels). For both evaluations, WAM outperforms EditGuard which, in particular, is not robust to cropping.
Figure 4: Evaluation of the localization on the validation set of COCO, with or without cropping before extraction, following the setup described in Sec. 5.4. (Left) Localization accuracy using intersection over union between the predicted watermarked areas and the ground-truth mask. (Right) Bit accuracy between the ground truth message and the decoded message, computed from Eq. (2).
5.5 Multiple watermarks
We compare the detection and decoding of multiple watermarks before and after the second training phase of WAM (Sec. 4.3). We embed up to five distinct messages into five separate 10% areas of every image (first resized to $256$ to ease the experiments). This is done by feeding the image several times to WAM’s embedder, then pasting the different watermarked areas onto the original image. These areas are squares disposed in a checkerboard pattern, which ensures that the watermarked areas do not overlap and have the same size (see Fig. 8 of App. 10.6). This is an arbitrary choice made to remove confounding factors during evaluation, although the training does not require the watermarked areas to be squared nor to have the same size. Following the methodology detailed in Sec. 3, we apply the DBSCAN algorithm to the vector outputs $\tilde{m}_{i}$ of the extractor corresponding to all pixels $i$ identified as watermarked ($y^{\text{det}}_{i}>\tau$). It identifies clusters corresponding to different watermarked regions without requiring the number of hidden messages to be known in advance. We use $\tau=0.5$ to threshold the watermarked pixels, and we choose a rather strict setup with $\varepsilon=1$ and $\text{min}_{\text{samples}}=1000$ pixels (around $2\%$ of the image) in our evaluations, resulting from an hyperparameter search detailed in Fig. 9 of App. 11.2. This means that a cluster will be considered only if it has at least 2% of the image’s pixels, and that the maximum distance between two predicted messages to be neighbors is 1.
Figure 5 presents the results. We compute the bit accuracies by comparing the centroid of each detected cluster (decoded message) with the ground truth message that has the largest overlapping area: the bit accuracy is thus computed only across the clusters that are discovered by DBSCAN. Without the second phase of training, the messages get mixed up, and WAM predicts the same (wrong) message for all watermarked pixels. After the training, WAM is able to accurately extract up to five different 32-bit messages from the areas covering 10% of the image each. Therefore, although the second phase of training embeds between 1 and 3 watermarks per image, WAM generalizes to more watermarks. It also shows that WAM’s effective capacity is greater than 32 bits (since it can hide multiple messages) if we do not consider heavy crops as a transformation we want the method to be robust against, similarly as (Bui et al., 2023b; Tancik et al., 2020; Wen et al., 2023).
This method remains effective after some image manipulations; extended quantitative results are presented in Fig. 9 of App. 11.2. For example, WAM achieves 85% mIoU even when the images have been horizontally flipped and the contrast adjusted. Furthermore, the bit accuracy, averaged across $5$ messages totaling 160 bits, exceeds 95% under the same conditions. However, WAM fails when JPEG compression is also added on top of these two transformations. Illustrative examples of the identified clusters in various scenarios can be found in Fig. 8 of App. 10.6.
(a) After first training phase (PSNR $\approx 25$ dB)
(b) After second training phase (PSNR $\approx 38$ dB)
Figure 5: Results on multiple watermarks extracted from a single image. We use non overlapping 10 % rectangular masks to watermark up to 5 parts of images from COCO, with different messages, and report the average number of clusters detected by DBSCAN, bit accuracy across found messages, as well as the mIoU of watermark detection on all objects. (Left) After the first training phase, the bit accuracy strongly decreases as the number of watermarks grows. (Right) After fine-tuning, it stays roughly constant no matter the number of watermarked parts. The mIoU stays stable in both cases.
6 Conclusion, Limitations, and Future Work
Conclusion.
This work introduces the Watermark Anything Model, which approaches image watermarking as a segmentation task. WAM is able to predict whether an image is watermarked or not, as well as to localize its watermarked regions. It thus handles images with a small watermarked part, or where parts of the image have been removed or edited. Additionally, it is able to detect and extract multiple watermarks within the same image, for the first time in the literature. Our training which introduces localization under heavy augmentations also offers state-of-the-art robustness on classical settings considered by current watermark methods, where most of the image is watermarked.
We identify two main limitations that we address in the next paragraphs.
Low payload.
In our experiments, WAM’s capacity is limited to 32 bits, and training on larger messages is challenging. In contrast, other watermarking methods such as EditGuard or RoSteALS can successfully embed more than 100 bits per image, but are not robust to crops or outpainting. This is a trade-off between capacity and robustness. Note that in practice, other methods use decoding to match with the original message and conclude if an image is watermarked. In contrast, WAM’s detection process is separate from the decoding, so 32 bits may be enough for most applications.
Perceptual quality.
In spite of the JND weighting, we notice that the watermark can still be visible in some areas of the watermarked images, even at a relatively high PSNR (see examples in App. 12). This could be due to the JND map focusing only on the cover image and not on the watermark signal itself. Improvements might be achieved by employing more sophisticated Human Visual System (HVS) models or by regularizing the watermark to eliminate repetitive patterns during training.
Reproducibility statement
Section 5 provides the specifics of our training and evaluation setup. Further implementation details, including network architectures, mask designs, transformation parameters for robustness evaluation, and settings for baseline comparisons, are presented in App. 10. The models and code for inference and training used in this paper will be made available upon publication. The code is based on PyTorch, the training dataset is publicly available, and the training requires half a week on 8 V100 GPUs, which makes the models reproducible at reasonable cost.
The training and inference code, as well as trained models are available at: https://github.com/facebookresearch/watermark-anything